# Meta-Generalized-Gradient Approximation made Magnetic

### v3

@@ Abstract
Introduction.  Despite nearly one hundred years of quantum physics, magnetism continues to challenge our understanding of electron correlations. An unsettling example is given by the recently developed strongly-constrained and appropriately normed (SCAN) meta-generalized gradient approximation (MGGA) for density functional theory (DFT).Sun et al. (2015, 2016)
@@
Here, $E_{\rm H}[n]$ and corresponding $v_{\rm H}$ are unchanged but the spin-dependent xc-potential, $v_{{\rm xc},\sigma}=\delta E_{\rm xc}/\delta n_{\sigma}$, is generated from the functional $E_{\rm xc}[n_{\uparrow},n_{\downarrow}]$, which depends both on $n_{\uparrow}$ and $n_{\uparrow}$ and not just on their sum $n=\sum_{\sigma}n_{\sigma}$.
@@
$E^{\rm nc}_{x}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{x}\left(n,\zeta_{\rm nc}\right)~{}F^{\rm SCAN}_{x}\left(n,s,\widetilde{\alpha}_{\rm nc}\right)\;,$ (13a)
@@
$E^{\rm nc}_{c}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta_{\rm nc})~{}F^{\rm SCAN}_{\rm c}(n,\zeta_{\rm nc},s,\widetilde{\alpha}_{\rm nc})\;,$ (13b)
@@
Notable magnetic SCAN failures — resolved. Let’s scrutinize the novel collinear solutions. We implemented the mSCAN functional in a developer’s version of the Crystal23 program.Erba et al. (2023) We study current-less, collinear magnets. Thus, the single-particle equation of SCDFT given in Eq. (10) is considerably simplified, as the vector potentials ${\bf A}_{\rm xc}$ and ${\bf A}^{a}_{\rm xc}$ vanish, and we must self-consistently solve the simpler Eq. (7) in a basis of Bloch orbitals $\varphi_{i,\sigma,\mathbf{k}}$ (with wavevector $\mathbf{k}$) formed from a linear combination of atom-centered spherical Gaussian-type functions (GTF).Desmarais et al. (2018) In Eq. (7) for mSCAN, in a generalized KS context,Seidl et al. (1996); Desmarais et al. (2024c) the spin-dependent xc potential $v_{{\rm xc},\sigma}$ readsNeumann et al. (1996)
$v_{{\rm xc},\sigma}=\left.\frac{\delta E^{\rm mSCAN}_{xc}}{\delta n^{\sigma}}\right|_{\tau^{\sigma}}-\nabla\cdot\frac{1}{2}\left.\frac{\delta E^{\rm mSCAN}_{xc}}{\delta\tau^{\sigma}}\right|_{n^{\sigma}}\cdot\nabla$ (16)
We solve Eqs. (7) and (16). Computational details are provided in the SM.

### v4

@@ Abstract
Introduction.  Nearly one hundred years after Heisenberg proposed the first quantum mechanical explanation of ferromagnetismHeisenberg (1928) magnetism continues to pose a challenge for ab initio electronic structure theory. An unsettling example is given by the recently developed strongly-constrained and appropriately normed (SCAN) meta-generalized gradient approximation (MGGA) for density functional theory (DFT).Sun et al. (2015, 2016)
@@
Here, $E_{\rm H}[n]$ and corresponding $v_{\rm H}$ are unchanged but the spin-dependent xc-potential, $v_{{\rm xc},\sigma}=\delta E_{\rm xc}/\delta n_{\sigma}$, is generated from the functional $E_{\rm xc}[n_{\uparrow},n_{\downarrow}]$, which depends both on $n_{\uparrow}$ and $n_{\downarrow}$ and not just on their sum $n=\sum_{\sigma}n_{\sigma}$.
@@
$E^{\rm nc}_{\rm x}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm x}\left(n,\zeta_{\rm nc}\right)~{}F^{\rm SCAN}_{\rm x}\left(n,s,\widetilde{\alpha}_{\rm nc}\right)\;,$ (13a)
@@
$E^{\rm nc}_{\rm c}[n,\vec{m},\mathbf{j},\vec{\mathbf{J}}]=\int d^{3}r~{}n~{}\varepsilon^{\rm unif}_{\rm c}(n,\zeta_{\rm nc})~{}F^{\rm SCAN}_{\rm c}(n,\zeta_{\rm nc},s,\widetilde{\alpha}_{\rm nc})\;,$ (13b)
@@
Notable magnetic SCAN failures — resolved. Let’s scrutinize the novel collinear solutions. We implemented the mSCAN functional in a developer’s version of the Crystal23 program.Erba et al. (2023) We study current-less, collinear magnets. Thus, the single-particle equation of SCDFT given in Eq. (10) is considerably simplified, as the vector potentials ${\bf A}_{\rm xc}$ and ${\bf A}^{a}_{\rm xc}$ vanish, and we must self-consistently solve the simpler Eq. (7) in a basis of Bloch orbitals $\varphi_{i,\sigma,\mathbf{k}}$ (with wavevector $\mathbf{k}$) formed from a linear combination of atom-centered spherical Gaussian-type functions (GTF).Desmarais et al. (2018) In Eq. (7) for mSCAN, in a generalized KS context,Seidl et al. (1996) the spin-dependent xc potential $v_{{\rm xc},\sigma}$ includes an additional semilocal term coming from the spin-kinetic energy dependence.Neumann et al. (1996); Lehtola et al. (2020); Desmarais et al. (2024c) Computational details for the general case are provided in the SM.